# Hestia: Hierarchical Next-Best-View Exploration for Systematic Intelligent Autonomous Data Collection

### v1

@@ I.2 Dataset Preparation
This section includes the details of the training and testing. We employ PPO [89] from stable-baselines3 [90] as the reinforcement learning (RL) framework for Hestia. The grid resolution $g$ is set to 20, and $h$ and $w$ are set to 300. The initial learning rate is $3\times 10^{-4}$ and is decayed by a factor of 2 every 500,000 iterations starting from 2,000,000 iterations until reaching 4,000,000, for a total of 5,000,000 training iterations. During training, $H_{t}$ is randomly sampled from the initial viewpoint height, up to a maximum of 10 meters. For testing, $H_{t}$ starts at 10 meters, is reduced to 5 meters during the last 10 to 5 steps, and further decreases to 2 meters in the final 5 steps. A training episode ends and the scene resets either when the number of captured images reaches 50 or when the target face coverage ratio of 0.9 is achieved. The complete training process takes approximately 24 hours on an NVIDIA RTX A6000 GPU.
@@
Our network architecture is lightweight, with only 4.9 million parameters (see Tab. A5), which is approximately half the size of a standard ResNet-18 model, which has around 11.7 million parameters. The proposal network consists of three 3D convolutional layers, each followed by a Leaky ReLU activation. This design progressively downsamples the 3D grid before further downstream operations. It is then followed by a 3D self-attention layer, again paired with a Leaky ReLU activation, to expand the receptive field. Finally, the network applies a reparameterization trick module, composed of linear layers, to generate the output distribution parameters. This design allows the look-at point to be sampled from a distribution rather than predicted deterministically. The grid encoder consists of three 3D convolutional layers, each followed by batch normalization and a Leaky ReLU activation. After encoding, trilinear interpolation is applied to each encoded grid feature, followed by feature concatenation. The image encoder is composed of three 2D convolutional layers, each followed by batch normalization and Leaky ReLU activation. The encoded features are then flattened and passed through a linear layer with Leaky ReLU activation. For the policy network, we adopt the default model provided in stable-baselines3. For more details about the network architecture, please refer to the code provided in the supplementary materials. The hierarchical design first predicts the look-at point, followed by the camera position. This design prioritizes the look-at point, as the primary objective in this task is to determine where to look rather than where to fly. It also resembles how a human pilot controls a drone during data capture, focusing first on the target of observation before planning the flight path.

### v2

@@ 6 Acknowledgements
This section includes the details of the training and testing. We employ PPO [schulman2017proximal] from stable-baselines3 [stable-baselines3] as the reinforcement learning framework for Hestia. The grid resolution $g$ is set to 20, and $h$ and $w$ are set to 300. The initial learning rate is $3\times 10^{-4}$ and is decayed by a factor of 2 every 500,000 iterations starting from 2,000,000 iterations until reaching 4,000,000, for a total of 5,000,000 training iterations. During training, $H_{t}$ is randomly sampled from the initial viewpoint height, up to a maximum of 10 meters. For testing, $H_{t}$ starts at 10 meters, is reduced to 5 meters during the last 10 to 5 steps, and further decreases to 2 meters in the final 5 steps. A training episode ends and the scene resets either when the number of captured images reaches 50 or when the target face coverage ratio of 0.9 is achieved. The complete training process takes approximately 24 hours on an NVIDIA RTX A6000 GPU.
@@
Our network architecture is lightweight, with only 4.9 million parameters (see Tab. 3, which is approximately half the size of a standard ResNet-18 model, which has around 11.7 million parameters. The proposal network consists of three 3D convolutional layers, each followed by a Leaky ReLU activation. This design progressively downsamples the 3D grid before further downstream operations. It is then followed by a 3D self-attention layer, again paired with a Leaky ReLU activation, to expand the receptive field. Finally, the network applies a reparameterization trick module, composed of linear layers, to generate the output distribution parameters. This design allows the look-at point to be sampled from a distribution rather than predicted deterministically. The grid encoder consists of three 3D convolutional layers, each followed by batch normalization and a Leaky ReLU activation. After encoding, trilinear interpolation is applied to each encoded grid feature, followed by feature concatenation. The image encoder is composed of three 2D convolutional layers, each followed by batch normalization and Leaky ReLU activation. The encoded features are then flattened and passed through a linear layer with Leaky ReLU activation. For the policy network, we adopt the default model provided in stable-baselines3. For more details about the network architecture, please refer to the code provided in the supplementary materials. The hierarchical design first predicts the look-at point, followed by the camera position. This design prioritizes the look-at point, as the primary objective in this task is to determine where to look rather than where to fly. It also resembles how a human pilot controls a drone during data capture, focusing first on the target of observation before planning the flight path.
S11 Limitations
Although the quantitative and qualitative results (see  Secs. 4.3 and S3) demonstrate a nearly comprehensive point cloud reconstruction, there are still some failure cases of Hestia (see Fig. S14). Hestia may occasionally fail to capture finer 3D structures, such as the window frames of the first-row house shown in Fig. S14. It may also fail to reconstruct highly self-occluded parts, such as the pillar of the second-row house. In addition, Hestia sometimes struggles to capture bottom-up views that require extreme vertical viewing angles, for example, the Lego man’s right hand and the underside of the pillar. Moreover, it may struggle to reconstruct shapes with fine details over coarse surfaces, such as tiny parts of the broccoli and anise. Adopting a multi-resolution grid structure or integrating octree-based methods to enhance the voxel grid resolution could be a potential future step to mitigate these issues.
In addition to the above limitations, we hope that Hestia will not be misused for other types of next-best-view (NBV) tasks. In this study, we found that a close-greedy training scheme can effectively mitigate spurious correlations and is well-suited to our problem definition (see Secs. S7, 3 and 4). However, the next-best-view problem is a broad research topic with varying objectives. This finding may not generalize to other NBV tasks, such as next-best-view for object tracking or next-best-view for human aesthetics, where long-term planning is more critical.
Due to hardware limitations (e.g., the absence of an RGB-D camera), Hestia cannot fully exhibit its potential in the real-world drone system. However, since we use a depth estimator to convert RGB images into depth maps, this limitation represents a trade-off rather than a fundamental constraint. Our experiments conducted in NVIDIA IsaacLab demonstrate the full capability of Hestia, while the real-world application highlights Hestia’s robustness when a depth sensor is unavailable. Furthermore, due to drone regulations, the real-world application of Hestia is conducted indoors using an indoor GPS system (e.g., HTC Lighthouse base stations). Drone policies vary across countries, and obtaining outdoor flight approvals can take up to a year in our region. Additionally, outdoor trials require significant funding, such as renting a safe test site measuring approximately 100 meters by 100 meters. As a future step, we plan to test Hestia outdoors to further validate its performance. Another future step is to extend Hestia to a multi-agent setting for large-scale outdoor scanning (e.g., city-scale) under power-constrained scenarios.